• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name miham

    Log entry time 00:03:17 on November 30, 2009

    Entry number 303605

    keyword=SWING Shift Summary (11/29/09)

    Owl shift summary: Miha Mihovilovic (SL),  Mark Dalton (TO)
                       John Arrington (3rd)
    16:00  Compton people are doing tests. We are not taking any data -
           waiting for them to finish. 
    16:26  Compton people gave up! :) They decided that we should go 
           back to production running. 
    16:33  Start beam recovery procedure. Requested for 2uA beam with 
           slits back to normal. Target is at EMPTY position.
    16:34  Beam is back. Beam position monitor readouts seem to be within 
           the range: 
           BPMA: x = -0.67, y = 0.082, 
           BPMB: x = -0.657, y = 0.120
    16:36  Ask MCC to do the HARP scan. They asked us to raise the current
           from 2uA to 5uA for the HARP scan - we agreed. 
    16:45  MCC calls and reports that they are having problems with the
           HARP IOC inside hallA. They need to secure beam in to our Hall
           and reboot IOC. 
    16:54  MCC reports that they are done with reboot and they want to send
           beam back to the hall.
    16:55  MCC wants to turn the raster OFF, to do the test. Target is 
    17:00  MCC finished the HARP scan. B monitor values are:  (x/y) = 
           180/140um and A monitor values are : (x/y) = 190/170um. 
           These values are within the limits. 
    17:06  RC decided to skip the BeO check and move directly to the LD2 
           target. We asked MCC to take beam away, mask motion so that we 
           can move the target. We also asked tMC to and change the raster 
           back to (4x4)MCC.
    17:15  Target at LD2 position. We asked MCC to give us 2uA with raster 
    17:22  Did the spot++ check (Run #25865). Spot looks fine. 
    17:24  Called MCC and asked them to increase the current to 110uA. We 
           are starting with the production. We are going to run only LEFT 
           arm and HAPPEX DAQs since HRS-R is down. 
    19:18  During Happex run #14322 ROC 25 crashed. We are trying to reboot 
           it but so far with no luck. At the moment we are not taking any 
    20:30  We received call from Bryan Hess from IT who reported, that he
           can  not reach a network switch in Left arm. We believe that ROC
           25 and network switch problems are related.
    20:41  We have problems with Compton ROC5 and ROC6. We were able to 
           reboot ROC5. To bring ROC 6 back,  we had to power cycle the 
           crate. Compton now works normally.
    20:52  We informed Bob M. about the Network switch problem, and he said 
           that we should check if other ROCs in left arm respond. They do 
           NOT. However, ROCs in right arm respond normally. Therefore we 
           decided to ask MCC for the escorted controlled access, so that 
           when  Bob arrives he will be able to go into the HALL and try to
           fix the problem.
    21:00  We are in controlled access.
    21:05  Bob et. al went into a HALL on theleft ARM. 
    21:30  Bob solved the problem with network switch. We rebooted all the 
           crates and restared both codas: HAPPEX and LEFT. Things seem to 
           work now. We are back in power permit.
    21:35  Asked MCC if they can give us our beam back: 110uA with (6x6)MCC
           raster. We will return directly to production, no beam recovery
    21:48  Called MCC and asked them to take HWP out.
    21:58  Started again with the production running. 
    22:15  CODA crashed. Now we have problems with right arm ROCs 26 and 28.
    22:20  We noticed that DAQ is not working again. Now ROC 26 and 28 do 
           not respond. Since symptoms were very similar to those in the 
           left arm an hour ago, we decided to call Bryan Hess from the IT 
           again, to check if the network switch in Right arm works. He 
           reported, that now he can not reach the switch in the right arm 
           and that this is very strange. Since this seems to be the same 
           problem as before but now in the Right arm we decided to call for
           the controlled access, go down into the hall,and try to power 
           cycle the switch in the right arm.      
    22:25  In controlled access.         
    22:30  Dianchen et. al went down the hall to fix the problem. He power 
           cycled the network switch.
    23:00  People came out of the hall. Coda and all the rocs seem to be 
           working again. We went back to beam permit, now waiting for the
    23:15  Beam is back. Start with the production. First good run is #25871.
    00:00  Shift ended during runs #14328 and  #25871.     
                                   RUN LIST
    Right Arm DAQ:
    Left Arm DAQ:
    25865: Spot++ check (A short run)
    25866: 17:25/18:20, I = 110uA, N = 3.7M, DT = 2%, Raster = (4x4)mm
    25867: JUNK
    25868: JUNK
    25869: JUNK
    25870: JUNK
    25871: 23:10/(passed 2 next shift), I = 110uA, DT = 3%, Raster=(4x4)mm 
    Happex DAQ:
    14321: 17:25/18:20, I = 110uA, N = 100k, DT = 2%, Raster = (4x4)mm
    14322: 18:20/19:18, I = 110uA, N = 100k, DT = 2%, Raster = (4x4)mm
           (ROC 25 and CODA chrashed during this run.) 
    14323: 19:18/?????, I = 110uA, N = 48k, DT = 2%, Raster = (4x4)mm   
           (ROC 25 and CODA chrashed during this run.) 
    14324: JUNK
    14325: JUNK
    14326: 21:50/22:10, I = 110uA, N = 20k, DT = 2%, Raster = (4x4)mm 
           (ROCs 26, 28 and CODA chrashed during this run.) 
    14327: JUNK
    14328: 23:10/(passed 2 next shift), I = 110uA, DT = 3%, Raster=(4x4)mm 

    A copy of this log entry has been emailed to: miham, sirca